When the dimensions of a cube are reduced by 4 inches on each side, the surface area of the cube is 864 square inches. What were the demensions of the original cube?
Answer by stanbon(75887)

When the dimensions of a cube are reduced by 4 inches on each side, the surface area of the cube is 864 square inches. What were the demensions of the original cube?
-----
Let the original dimensions be x*x*x
----
With the new dimensions you have:
(x-4)^3 = 864
----
x-4 = cbrt(864) = 2*cbrt(108)
----
Dimension of original cube: x = [4+2cbrt(108)] inches
==============
Cheers,
Stan H.
====================